Aplane flying 3458-mi from new york city to london has a 50mph tailwind. The flights point of no return is the point at which the flight time required to return to new york is the same as the time required to continue to london. If the planes speed in still air is 860 mph, how far is new york from the point of no return?
The distance from new york to the point of no return is ____mi.
(round to the nearest integer.)

Answer by mananth(16946)

Aplane flying 3458-mi from new york city to london has a 50mph tailwind. The flights point of no return is the point at which the flight time required to return to new york is the same as the time required to continue to london. If the planes speed in still air is 860 mph, how far is new york from the point of no return?
The distance from new york to the point of no return is ____mi.

still air speed = 860 mph
wind speed = 50 mph
t=d/r
let x be the distance of point of no return from NY
distance from London = 3458-x
time to London d/r= 3458-x/(860+50)
=(3458-x)/910
time to NY *against wind)
=x/(860-50)
=x/810
These times are same
(3458-x)/910=x/810
910x=810*(3458-x)
910x=2800980-810x
1720x=2800980
x=1628
the distance of point of no return from NY= 1628 miles